# Learning Performance Maximizing Ensembles with Explainability Guarantees

### 20 Dec 2023 v1

@@ Abstract
In this paper we propose a method for the optimal allocation of observations between an intrinsically explainable glass box model and a black box model. An optimal allocation being defined as one which, for any given explainability level (i.e. the proportion of observations for which the explainable model is the prediction function), maximizes the performance of the ensemble on the underlying task, and maximizes performance of the explainable model on the observations allocated to it, subject to the maximal ensemble performance condition. The proposed method is shown to produce such explainability optimal allocations on a benchmark suite of tabular datasets across a variety of explainable and black box model types. These learned allocations are found to consistently maintain ensemble performance at very high explainability levels (explaining $74\%$ of observations on average), and in some cases even outperforming both the component explainable and black box models while improving explainability.
@@ Introduction
Explainability, however defined (Linardatos, Papastefanopoulos, and Kotsiantis 2020), is a desireable characteristic in any prediction function. Intrinsically interpretable “glass box” models ((Agarwal et al. 2021), (Lemhadri, Ruan, and Tibshirani 2021), (Rymarczyk et al. 2020)), which are explainable by construction, are particularly advantageous as they require no additional post-hoc processing ((Ribeiro, Singh, and Guestrin 2016), (Lundberg and Lee 2017)) to achieve explainability, and thus also avoid complications arising from post-hoc explanation learning ((Rudin 2019), (Garreau and Luxburg 2020)). Due to these advantages, glass box models are uniquely suited to settings where faithful explanations of predictions are required.
@@
Figure 1: This figure shows a two class classification task in which the areas of expertise (the diamond pattern for the glass box and the spiral pattern for the black box model) are complementary. In this example, the glass box achieves a $92.7\%$ accuracy, the black box reaches $95.0\%$ accuracy, and the allocated ensemble of the two exceeds both with a $95.8\%$ accuracy. Thus, the resulting EEG allocation improves performance over both component models while also providing explainability (for $20\%$ of observations in this case).
@@
Finally, methods are characterized by their allocation criteria which commonly consist of an objective which combines one or more of the following - the explainability level, the underlying task performance of the ensemble, and the complexity of the glass box model. Most methods optimize a measure of post-allocation ensemble performance - LSP, HyRS, HyPM, and HybridCORELS minimize a 0/1 misclassification loss, AdaBudg uses a more flexible logistic loss, and CRL maximizes accuracy across a range of explainability levels. Several methods with rule-based glass box/allocator hybrid models (HyRS, HyPM, CRL, and HybridCORELS) also include a penalty on the complexity of these models. To control the explainability level, methods either include a reward term in the loss (HyRS, HyPM, and CRL), or directly restrict the model space to candidates which achieve the explainability level (HybridCORELS). In contrast, EEG optimizes an MSE loss between the predicted and actual “glass box allocation desireability” percentile of each observation.
@@
Allocation Desireability Ranking: The second novel aspect of the EEG approach is the concept of allocation desireability. Given an ensemble of models, allocation desireability quantifies how beneficial it is for a given observation to be allocated to the default ensemble member model, say the glass box. Thus, it induces a preference for glass box allocation between all pairs of observations and consequently also defines a ranking of allocation preference across all observations that is optimal irrespective of the desired explainability level.
A key advantage of such a ranking is that it is independent of the training criteria of the ensemble member models, and thus can be adapted to score allocation desireability using metrics that best fit the setting. Indeed, the EEG desireability metric builds a ranking using a combination of relative sufficient performance and absolute performance measures which can natively accommodate any underlying problem type (e.g. regression, classification). This particular desireability metric also offers several additional benefits including allocation desireability percentile and sufficiency category estimates for each observation.
Q-Complete Allocation Optimality: The final key point of novelty of the EEG approach is the optimality of allocation, as defined in Proposition 1 and Proposition 2, which is encoded in the allocation desireability ranking for any explainability level. Thus, the learned allocator, which estimates this ranking, is an explicit function of $q$ and provides the allocation solution to any explainability level after training only once. This capability is in contrast with previous works which provide, at most, several explainability level solutions with varying degrees of stability (Ferry, Laberge, and Aïvodji 2023).
@@
  • Given a minimum performance requirement on the underlying task, the method can be used to obtain the allocation with the highest interpretability level that achieves or exceeds the performance threshold.
@@
  • Given a minimal level of post-allocation glass box-specific performance, the allocation that achieves the highest interpretability level while meeting or exceeding this requirement can be found.
@@
In this case, loss minimization dictates an allocation of $z_{1}$ to $g$ and $z_{2}$ to $b$, which would allocate $z_{2}$ to an insufficient prediction. Sufficiency maximization would however yield a more satisfactory allocation of $z_{2}$ to $g$ and $z_{1}$ to $b$. This example demonstrates the utility of sufficiency allocation - distinguishing between a case where the user is willing to sacrifice “a bit of performance” (as quantified by sufficiency) for explanation ($z_{1}$), and a case were even a small drop performance results in an explanation is not sufficiently trustworthy to use ($z_{2}$).
@@
The intuition behind the allocator is as follows. First, all observations are sorted in sufficient performance maximizing order, i.e. allocation of groups $Z_{g}^{\prime}$ and $Z_{2}^{\prime}$ to $g$ is prioritized over allocation of groups $Z_{0}^{\prime}$ and $Z_{b}^{\prime}$ to $g$. Next, within each group, observations are sorted in explainable sufficient performance maximizing order, i.e. $Z_{g}^{\prime}$ is prioritized ahead of $Z_{2}^{\prime}$ and $Z_{0}^{\prime}$ is prioritized ahead of $Z_{b}^{\prime}$ for allocation to $g$. Then, within each sufficiency category, observations are ordered in absolute performance maximizing order, i.e. observations with large relative performance of $g$ over $b$ are prioritized for allocation to $g$. Next, this ranking is normalized, yielding the glass box allocation desireability percentile $r^{\prime}$. An important feature of this percentile is that it is constant with respect to $q$, thus the optimal observations to allocate to $g$, for any level of $q$, are simply the $n_{q}$ most highly ranked, resulting in the allocator $a^{\prime}_{q}$.
@@
Next, we state the optimality properties of the proposed allocator $a^{\prime}_{q}$ (with proofs available in the Theoretical Results section of the Appendix).
@@
Each glass box and black box model is tuned on the full set of hyperparamters each time it is replicated. The gradient boosting trees allocator models are retuned on the full hyperparameter set each time as well. The neural network allocator is not retuned however, and instead uses the optimal settings found in the fitting of the black box on each dataset.
@@
Next, we define the $95\%$ Threshold Q Max (95TQM) as the highest value of $q$ for which the ensemble performance meets or exceeds $95\%$ of the performance of the better of $g$ and $b$. Thus this is a measure of how much interpretability can be utilized before the performance price becomes material.
Next, we define the maximum accuracy achieved by the allocator across all $q$ (Max Acc), and the highest value of $q$ for which this accuracy is maintained (Argmax $q$). The Max Acc can be benchmarked against the AUC, interpretable as the average accuracy across $q$. Each of these metrics is a percentage and higher values correspond with higher performance and higher interpretability at this maximum performance level, respectively.
@@ Results
Evaluation of allocator performance using the metrics defined previously as well as visual inspection of the performance vs interpretability trade-off curves (Fig. 2) revealed both the benefits and some of the limitations of learned allocation in the tabular data setting.
Firstly, performance was found to consistently and significantly outperform random allocation, as quantified by a cross-dataset PPCR of $0.37\pm 0.21$ (Table 1), indicating that the learned allocation captured close to $40\%$ of the area under the curve available and in excess of random allocation. It was also found that on some datasets in particular, learned allocation performed close to oracle allocation (e.g. $89\%$ and $71\%$ on the Isolet and BrazilianHouses datasets).
Learned allocation was also found to perform at least at the level of the best ensemble member model across an average of $74\%$ of the explainability range (PQEOM in Table 1). This indicates that for many datasets, there is a substantial explainability “free lunch” to be taken advantage of without performance loss. On a few datasets, performance of the allocated ensemble was found to outperform both $g$ and $b$ for a majority ($93\%$) of the $q$ range (PolR and FifaR PQOM). The 95TQM also supported these conclusions, with a cross-dataset average value of $0.94\%\pm 0.09$ indicating that allocation performance was within $5\%$ of maximal indiviudal model performance across approximately all values of $q$.
Assessing the PCFA metric suggests some limits to the upside of learned, feature dependent allocation - at least in the tested tabular data setting. A cross-dataset average value of $0.35\pm 0.34$ indicates that on average, the range for which the feature dependent allocator is used over the feature-independent one is indistinguishable from zero. This is consistent with a visual inspection of the representative performance-interpretability curve e.g. Fig. 2 (b) where there is no improvement to be had in excess of $a^{\prime\prime}_{q}$. However, it is noted that the only possibility for “homerun” allocations is through the feature dependent $a^{\prime}_{q}$ as seen in Fig 2 (a) with the PolR dataset and also in Table 1 for datasets SulfurR, BikeSharingR, and FifaR. Thus the ensembled allocation scheme offers this upside without downside risk of low performance in either $a^{\prime}_{q}$ or $a^{\prime\prime}_{q}$.
Evaluation of the case in which one single allocation is needed is also positive. On a cross-dataset average, the $84\%$ maximal accuracy achieved is quite high, and is also achieved at a high average explainability level ($64\%$). Particularly strong individual results can be seen in the Pol and SulfurR datasets (Table 1). We also find that on a observation-level, the allocation is an accurate estimator of sufficiency category, with a cross-dataset average of $76\%$ and with few datasets with accuracy under $60\%$.
@@
Figure 2: This figure shows two examples of the interpretability vs performance trade-off, comparing the random, oracle, and learned allocation curves. The PolR dataset is an example of complementary $g$ and $b$ models, resulting in an allocated ensemble that outperforms both component models across most of the $q$ range. The SuperconductR dataset is an example of an explainability “free lunch” in which the $b$ accuracy is maintains while increasing explainability using the allocator. Curves for all datasets available in Fig. 3 in the Appendix.
@@ Appendix A Appendix
Figure 3: This figure provides the sufficient accuracy values of the allocated EEG ensemble for each explainability level and every dataset.

### 22 Dec 2023 v2

@@ Abstract
In this paper we propose a method for the optimal allocation of observations between an intrinsically explainable glass box model and a black box model. An optimal allocation being defined as one which, for any given explainability level (i.e. the proportion of observations for which the explainable model is the prediction function), maximizes the performance of the ensemble on the underlying task, and maximizes performance of the explainable model on the observations allocated to it, subject to the maximal ensemble performance condition. The proposed method is shown to produce such explainability optimal allocations on a benchmark suite of tabular datasets across a variety of explainable and black box model types. These learned allocations are found to consistently maintain ensemble performance at very high explainability levels (explaining $74\%$ of observations on average), and in some cases even outperform both the component explainable and black box models while improving explainability.
@@ Introduction
Explainability, however defined (Linardatos, Papastefanopoulos, and Kotsiantis 2020), is a desirable characteristic in any prediction function. Intrinsically interpretable “glass box” models ((Agarwal et al. 2021), (Lemhadri, Ruan, and Tibshirani 2021), (Rymarczyk et al. 2020)), which are explainable by construction, are particularly advantageous as they require no additional post-hoc processing ((Ribeiro, Singh, and Guestrin 2016), (Lundberg and Lee 2017)) to achieve explainability, and thus also avoid complications arising from post-hoc explanation learning ((Rudin 2019), (Garreau and Luxburg 2020)). Due to these advantages, glass box models are uniquely suited to settings where faithful explanations of predictions are required.
@@
Figure 1: This figure shows a two-class classification task in which the areas of expertise (the diamond pattern for the glass box and the spiral pattern for the black box model) are complementary. In this example, the glass box achieves a $92.7\%$ accuracy, the black box reaches $95.0\%$ accuracy, and the allocated ensemble of the two exceeds both with a $95.8\%$ accuracy. Thus, the resulting EEG allocation improves performance over both component models while also providing explainability (for $20\%$ of observations in this case).
@@
Finally, methods are characterized by their allocation criteria which commonly consist of an objective which combines one or more of the following - the explainability level, the underlying task performance of the ensemble, and the complexity of the glass box model. Most methods optimize a measure of post-allocation ensemble performance - LSP, HyRS, HyPM, and HybridCORELS minimize a 0/1 misclassification loss, AdaBudg uses a more flexible logistic loss, and CRL maximizes accuracy across a range of explainability levels. Several methods with rule-based glass box/allocator hybrid models (HyRS, HyPM, CRL, and HybridCORELS) also include a penalty on the complexity of these models. To control the explainability level, methods either include a reward term in the loss (HyRS, HyPM, and CRL), or directly restrict the model space to candidates which achieve the explainability level (HybridCORELS). In contrast, EEG optimizes an MSE loss between the predicted and actual “glass box allocation desirability” percentile of each observation.
@@
Allocation Desirability Ranking: The second novel aspect of the EEG approach is the concept of allocation desirability. Given an ensemble of models, allocation desirability quantifies how beneficial it is for a given observation to be allocated to the default ensemble member model, say the glass box. Thus, it induces a preference for glass box allocation between all pairs of observations and consequently also defines a ranking of allocation preference across all observations that is optimal irrespective of the desired explainability level.
A key advantage of such a ranking is that it is independent of the training criteria of the ensemble member models, and thus can be adapted to score allocation desirability using metrics that best fit the setting. Indeed, the EEG desirability metric builds a ranking using a combination of relative sufficient performance and absolute performance measures which can natively accommodate any underlying problem type (e.g. regression, classification). This particular desirability metric also offers several additional benefits including allocation desirability percentile and sufficiency category estimates for each observation.
Q-Complete Allocation Optimality: The final key point of novelty of the EEG approach is the optimality of allocation, as defined in Proposition 1 and Proposition 2, which is encoded in the allocation desirability ranking for any explainability level. Thus, the learned allocator, which estimates this ranking, is an explicit function of $q$ and provides the allocation solution to any explainability level after training only once. This capability is in contrast with previous works which provide, at most, several explainability level solutions with varying degrees of stability (Ferry, Laberge, and Aïvodji 2023).
@@
  • Given a minimum performance requirement on the underlying task, the method can be used to obtain the allocation with the highest explainability level that achieves or exceeds the performance threshold.
@@
  • Given a minimal level of post-allocation glass box-specific performance, the allocation that achieves the highest explainability level while meeting or exceeding this requirement can be found.
@@
In this case, loss minimization dictates an allocation of $z_{1}$ to $g$ and $z_{2}$ to $b$, which would allocate $z_{2}$ to an insufficient prediction. Sufficiency maximization would however yield a more satisfactory allocation of $z_{2}$ to $g$ and $z_{1}$ to $b$. This example demonstrates the utility of sufficiency allocation - distinguishing between a case where the user is willing to sacrifice “a bit of performance” (as quantified by sufficiency) for explanation ($z_{1}$), and a case were even a small performance drop results in an explanation that is not sufficiently trustworthy to use ($z_{2}$).
@@
The intuition behind the allocator is as follows. First, all observations are sorted in sufficient performance maximizing order, i.e. allocation of observations in $Z_{g}^{\prime}$ to $g$ is prioritized over allocation of observations in $Z_{2}^{\prime}$ and $Z_{0}^{\prime}$, which in turn are prioritized over $Z_{b}^{\prime}$. Next, observations are sorted in explainable sufficient performance maximizing order, i.e. $Z_{2}^{\prime}$ is prioritized ahead of $Z_{0}^{\prime}$ for allocation to $g$. Then, within each sufficiency category, observations are ordered in absolute performance maximizing order, i.e. observations with large relative performance of $g$ over $b$ are prioritized for allocation to $g$. Next, this ranking is normalized, yielding the glass box allocation desirability percentile $r^{\prime}$. An important feature of this percentile is that it is constant with respect to $q$, thus the optimal observations to allocate to $g$, for any level of $q$, are simply the $n_{q}$ most highly ranked, resulting in the allocator $a^{\prime}_{q}$.
@@
Next, we state the optimality properties of the proposed allocator $a^{\prime}_{q}$. The proofs are available in the long form paper on arxiv.org in the Theoretical Results section of the Appendix.
@@
Each glass box and black box model is tuned on the full set of hyperparameters each time it is replicated. The gradient boosting trees allocator models are retuned on the full hyperparameter set each time as well. The neural network allocator is not retuned however, and instead uses the optimal settings found in the fitting of the black box on each dataset.
@@
Next, we define the $95\%$ Threshold Q Max (95TQM) as the highest value of $q$ for which the ensemble performance meets or exceeds $95\%$ of the performance of the better of $g$ and $b$. Thus this is a measure of how much explainability can be utilized before the performance price becomes material.
Next, we define the maximum accuracy achieved by the allocator across all $q$ (Max Acc), and the highest value of $q$ for which this accuracy is maintained (Argmax $q$). The Max Acc can be benchmarked against the AUC, interpretable as the average accuracy across $q$. Each of these metrics is a percentage and higher values correspond with higher performance and higher explainability at this maximum performance level, respectively.
@@ Results
Evaluation of allocator performance using the metrics defined previously as well as visual inspection of the performance vs explainability trade-off curves (Fig. 2) revealed both the benefits and some of the limitations of learned allocation in the tabular data setting.
Firstly, performance was found to consistently and significantly outperform random allocation, as quantified by a cross-dataset PPCR of $37\%$ (Table 1), indicating that the learned allocation captured close to $40\%$ of the area under the curve available and in excess of random allocation. It was also found that on some datasets in particular, learned allocation performed close to oracle allocation (e.g. $89\%$ and $71\%$ on the IsoletR and BrazilianHousesR datasets).
Learned allocation was also found to perform at least at the level of the best ensemble member model across an average of $74\%$ of the explainability range (PQEOM in Table 1). This indicates that for many datasets, there is a substantial explainability “free lunch” to be taken advantage of without performance loss. On a few datasets, performance of the allocated ensemble was found to outperform both $g$ and $b$ for a majority ($93\%$) of the $q$ range (PolR and FifaR PQOM). The 95TQM metric also supported these conclusions, with a cross-dataset average value of $94\%$ indicating that allocation performance was within $5\%$ of maximal individual model performance across approximately all values of $q$.
Assessing the PCFA metric suggests some limits to the upside of learned, feature dependent allocation - at least in the tested tabular data setting. A cross-dataset average value of $35\%\pm 34\%$ indicates that on average, the range for which the feature dependent allocator is used over the feature-independent one is indistinguishable from zero. This is consistent with a visual inspection of the representative performance-explainability curve e.g. Fig. 2 (b) where there is no improvement to be had in excess of $a^{\prime\prime}_{q}$. However, it is noted that the only possibility for “homerun” allocations is through the feature dependent $a^{\prime}_{q}$ as seen in Fig 2 (a) with the PolR dataset and also in Table 1 for datasets SulfurR, BikeSharingR, and FifaR. Thus the ensembled allocation scheme offers this upside without downside risk of low performance in either $a^{\prime}_{q}$ or $a^{\prime\prime}_{q}$.
Evaluation of the case in which a single allocation is needed is also positive. On a cross-dataset average, the $84\%$ maximal accuracy achieved is quite high, and is also achieved at a high average explainability level ($64\%$). Particularly strong individual results can be seen in the Pol and SulfurR datasets (Table 1). We also find that on a observation-level, the allocation is an accurate estimator of sufficiency category, with a cross-dataset average of $76\%$ and with few datasets with accuracy under $60\%$.
@@
Figure 2: This figure shows two examples of the explainability vs performance trade-off, comparing the random, oracle, and learned allocation curves. The PolR dataset is an example of complementary $g$ and $b$ models, resulting in an allocated ensemble that outperforms both component models across most of the $q$ range. The SuperconductR dataset is an example of an explainability “free lunch” in which the $b$ accuracy is maintained while increasing explainability using the allocator. Curves for all datasets are available in the Appendix of the long form paper available on arxiv.org.
@@ Appendix A Appendix
Figure 3: This figure provides the sufficient accuracy values of the allocated EEG ensemble for each explainability level and every dataset using the neural network allocator.
Figure 4: This figure provides the sufficient accuracy values of the allocated EEG ensemble for each explainability level and every dataset using the gradient boosting trees allocator.